OpenStudy (anonymous):

A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per second. What is the maximum height of the particle?

OpenStudy (anonymous):

helpme

OpenStudy (anonymous):

Assuming Earth's gravity, the formula for the flight of the particle is: s(t) = -16t^2 + vt + s = -16t^2 + 144t + 160. This has a maximum when t = -b/(2a) = -144/[2(-16)] = -144/(-32) = 9/2. Hence , the maximum height is s(9/2) = -16(9/2)^2 + 144(9/2) + 160 = 484 feet.

OpenStudy (anonymous):

your a lifesaver

OpenStudy (anonymous):

haha thank you :)